THZ SPECTROSCOPY OF ^{13}C ISOTOPIC SPECIES OF A ”WEED”: ACETALDEHYDE

\underline{L. MARGULÈS}, and R. A. MOTIYENKO, Laboratoire PhLAM, CNRS UMR 8523, Université de Lille 1, 59655 Villeneuve d’Ascq Cedex, France; and J.-C. GUILLEMIN, Sciences Chimiques de Rennes, UMR 6226 CNRS-ENSCR, Avenue du Général Leclerc, CS 50837, 35708 Rennes Cedex 7, France.

Our studies of the isotopic species of ^{13}C and D isotopologues of methyl formate (HCOOCH_{3}), have allowed the detection of more than 600 lines in Orion^{a}^{,}^{b}. This confirms that many observed U-lines are coming from isotopic species of one of the most abundant molecules in space. Since its first detection in 1976 in SgrB2 and in Orion A, acetaldehyde (CH_{3}CHO) was detected in many other numerous objects^{c}. If its deuterated species (CD_{3}CHO^{d} and CH_{3}CDO^{e}) have been previously studied in the millimeterwave range, the data concerning the ^{13}C species are limited to few lines measured in 1957 up to 40 GHz^{f}. In this context we decided to study the ^{13}C species of acetaldehyde. Acetaldehyde molecule displays a large amplitude motion: the hindered rotation of the methyl group with respect to the rest of the molecule. The analysis is performed with the Rho Axis Method^{g}. Recent versions of the codes include high orders term in order to reproduce the observed frequencies for large quantum numbers values as J-values as high as 70^{a,b,}^{h}. Measurements and analysis of the rotational spectra of ^{13}C isotopic species are in progress in Lille with a solidstate submillimetre-wave spectrometer (50-950 GHz), the first results will be presented. This work is supported by the contract ANR-08-BLAN-0054 and by the Programme National de Physico-Chimie du Milieu Interstellaire (PCMI-CNRS).

aCarvajal, M.; Margulès, L.; Tercero, B.; et al.A&A 500, (2009) 1109

bMargulès, L.; Huet, T. R.; Demaison J.; et al.,ApJ 714, (2010) 1120

cIkeda, M.; Ohishi, M.; Nummelin, A.; et al., ApJ, 560, (2001) 792

dKleiner, I.; Lopez, J.-C.; Blanco, S.; et al.J. Mol. Spectrosc. 197, (1999) 275

eElkeurti M.; Coudert, L. H.; Medvedev, I. R.; et al.J. Mol. Spectrosc. 263, (2010) 145

fKilb, R.W.; Lin, C.C.; and Wilson, E.B.J. Chem. Phys. 26, (1957) 1695

gKleiner, I. J. Mol. Spectrosc. 260, (2010) 1

hIlyushin, V.V.; Kryvda, A; and Alekseev, E;J. Mol. Spectrosc. 255, (2009) 32